OpenStudy (anonymous):

find the distance between the following points: (-1,-7) (4,5)

OpenStudy (akashdeepdeb):

Use the distance formula! :) Do you know it?

OpenStudy (anonymous):

I'm looking for it??

OpenStudy (akashdeepdeb):

Distance between 2 points = \[\sqrt{(x2-x1)^2 + (y2-y1)^2}\] Use the points given in the equation! :)

OpenStudy (jdoe0001):

\(\bf \text{distance between 2 points}\\ d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\)

OpenStudy (anonymous):

=169

OpenStudy (akashdeepdeb):

Yes excellent! But the square root of that right? It is under root! :D

OpenStudy (akashdeepdeb):

@Algebra2013 ? :D

OpenStudy (anonymous):

=169? or \[169^2\] ??

OpenStudy (akashdeepdeb):

No see You must have got it as sqrt 169 right?

OpenStudy (anonymous):

no just 169

OpenStudy (akashdeepdeb):

So \[\sqrt{169}\] = +13 or -13 And we know it is distance so it cannot be negative.... So the distance between them would be 13 units!! :D Understood? :)

OpenStudy (anonymous):

oh ok. cool!!!!!!!!!! so what would 162 be?? as a square

OpenStudy (anonymous):

\[9\sqrt{2}\]

OpenStudy (akashdeepdeb):

For sqrt 162 Right it in terms of it factors!! :D \[\sqrt{162} = \sqrt{2 * 9 * 9} = 9\sqrt{2}\] Got it? :) YES CORRECT!! :D

OpenStudy (akashdeepdeb):

Good job! :)